A thirty year monthly payment mortgage loan for 250,000 is offered at a nominal rate of 6.3% convertible monthly. The lender charges a fee of 2.5% for which no services are provided. Find the APR.
Answer #1

Answer -

An APR or Annual Percentage Rate is the real rate of interest of a loan after including/incorporating all associated costs involved. It is generally higher than the rate of interest offered in the loan documentation since other incidental costs like handling fees, processing fees, documentation fees etc are always extra.

Formula for calculating APR is (All fees + Interest amount)/Principal * 100.  

In our case, the following are the factual data

Loan amount = 250000

Interest Rate = 6.3%

Fees = 2.5%

Tenor = 30 years

Hence the interest amount is 250000*6.3% = 15750. Fees amount is 250000*2.5% = 6250. The total outflow is 15750+6250 = 22000. This total outflow as a percentage of the principal amount would give you the APR.

APR = 22000/250000 = 8.8%.

APR is 8.8% for the loan.